Find the indicated limit, if it exists.

limit of f of x as x approaches 0 where f of x equals 10 x plus 2 when x is less than 0 and the absolute value of the quantity 2 minus x when x is greater than or equal to 0
1
Expert's answer
2019-08-21T04:10:44-0400

Answer

The limit exists and equals 2.

Explanation

Since f(x)=10x+2 for x"\isin ]-\infin; 0[" and it is continuous on ]"-\infin"; 0[ as a polynomial, then its left limit at the point x=0 exists and equals f(0-0)=f(0)=10*0+2=2.

On the other hand, f(x)=|2-x| for x "\isin [0,+\infin[" and it is continuous on [0,+"\infin"[ as the absolute value of a polynomial, then the right limit at the point x=0 equals

f(0+0)= f(0)=|2-0|=2.

Hence, f(0-0)=f(0+0)=2, and the limit of f (x) as x approaches 0 exists and equals 2.